OpenStudy (anonymous):

What are the square roots of 196? A. 13 and –13 B. 14 and –14 C. 16 and –16 D. 98 and –98

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: we have: 14^2=196, and (-14)*(-14)= 196

OpenStudy (anonymous):

that's what i got. how about another question?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

What are the square roots of 576? A. −14 and 14 B. −24 and 24 C. −28 and 28 D. −288 and 288

OpenStudy (michele_laino):

hint: 576=24*24 and: 576=(-24)*(-24)

OpenStudy (anonymous):

that's what i got. next question

OpenStudy (anonymous):

What are the square roots of 64 over 144? A. minus 4 over 12 and 4 over 12 B. minus 4 over 9 and 4 over 9 C. minus 8 over 12 and 8 over 12 D. minus 32 over 72 and 32 over 72

OpenStudy (michele_laino):

hint: 64=8*8 144=12*12 so: \[\sqrt {\frac{{64}}{{144}}} = ...?\]

OpenStudy (anonymous):

when i divided 64 by 144 i got 0.44444444444444444444444444444444

OpenStudy (michele_laino):

hint: \[\sqrt {\frac{{64}}{{144}}} = \pm \sqrt {\frac{{{8^2}}}{{{{12}^2}}}} = ...?\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 9 } and -\frac{ 4 }{ 9 } \]

OpenStudy (michele_laino):

no, since: \[\begin{gathered} \sqrt {64} = \pm 8 \hfill \\ \sqrt {144} = \pm 12 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

when i divided 64 by 144 i got 0.44444444444444444444444444444444. so i went through each answer choice and divided the numerator by the denominator. When i divided 4 by 9 i got "0.44444444444444444444444444444444." so wouldn't that mean that 4 over 9 and negative 4 over 9 is the answer?

OpenStudy (michele_laino):

Please, I have to go to have my dinner, I will answer to your questions as soon as possible

OpenStudy (anonymous):

ok